OpenStudy (angelicahicks29):

sec^2(x)-1=csc^2(x)-cot^2(x)+sin^2(x)

OpenStudy (jdoe0001):

so hmmm is that a question or a statement?

OpenStudy (angelicahicks29):

A question..

OpenStudy (jdoe0001):

seems a bit occult though what would the question be though?

OpenStudy (angelicahicks29):

I have to make these equal..

OpenStudy (angelicahicks29):

This is a problem in my math book

OpenStudy (jdoe0001):

hmmm identity verification...h....k

OpenStudy (angelicahicks29):

\[\sec ^{2}x-1=\csc^2x-\cot^2x+\sin^2x\]

OpenStudy (angelicahicks29):

It looks better that way

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

let us start by noticing something \(\bf \textit{Pythagorean Identities} \\ \quad \\ sin^2(\theta)+cos^2(\theta)=1 \\ \quad \\ 1+cot^2(\theta)=csc^2(\theta) \\ \quad \\ 1+tan^2(\theta)=sec^2(\theta)\implies tan^2(\theta)={\color{brown}{ sec^2(\theta)-1 }} \)

OpenStudy (jdoe0001):

so we could say the left side is really just \(tan^2(x)\)

OpenStudy (angelicahicks29):

I noticed I wrote the wrong side down after rewriting everything. The real equation is \[1+\sin^2x=\csc^2x-\cot^2x+\sin^2x\]

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

alright let us poke the right-hand side then

OpenStudy (angelicahicks29):

since the right hand side is an identity can't we just cancel those?

OpenStudy (jdoe0001):

hmmm actaully lemme fix that quick, I have the old version

OpenStudy (jdoe0001):

\(\bf 1+sin^2(x)=csc^2(x)-cot^2(x)+sin^2(x) \\ \quad \\ \quad \\ \\ \quad \\ csc^2(x)-cot^2(x)+sin^2(x)\implies \cfrac{1}{sin^2(x)}-\cfrac{cos^2(x)}{sin^2(x)}+\cfrac{sin^2(x)}{1} \\ \quad \\ \cfrac{1-cos^2(x)+sin^2(x)sin^2(x)}{sin^2(x)} \\ \quad \\ {\color{brown}{ sin^2(\theta)+cos^2(\theta)=1\implies sin^2(\theta)=1-cos^2(\theta) }}\qquad thus \\ \quad \\ \cfrac{1-cos^2(x)+sin^2(x)sin^2(x)}{sin^2(x)}\implies \cfrac{{\color{brown}{ sin^2(x)}}+sin^2(x)sin^2(x)}{sin^2(x)} \\ \quad \\ \cfrac{\cancel{sin^2(x)}[1+sin^2(x)]}{\cancel{sin^2(x)}}\impliedby \textit{common factor}\)

OpenStudy (jdoe0001):

but yes, you're correct you can just use the \(\bf 1+cot^2(\theta)=csc^2(\theta)\implies 1=csc^2(\theta)-cot^2(\theta)\) and not cancel out, but use the "1"

OpenStudy (angelicahicks29):

then it becomes \[1+\sin^2x\]

OpenStudy (jdoe0001):

yeap

OpenStudy (angelicahicks29):

Another question!!

OpenStudy (angelicahicks29):

factoring \[\sin^3x-\cos^3x\]